OpenStudy (anonymous):

How do I find the length of the major and minor axis of (x-5)^2/4 + (y+3)^2/16=1?

OpenStudy (anonymous):

I got that the center is (5,-3) Vertices = (5,1) (5,-7) Foci = (5,-3+2√3) (5,-3-2√3) but I cannot find the length of the major and minor axis or the eccentricity...

OpenStudy (luigi0210):

Do you know what they are?

OpenStudy (anonymous):

no, I don't :(

OpenStudy (luigi0210):

Well if I'm not mistaken, this is an ellipse right?

OpenStudy (luigi0210):

So that means for the axis: Major axis: The longest diameter of an ellipse. Minor axis: The shortest diameter of an ellipse.

OpenStudy (anonymous):

yes, you are correct! but how do I find what the major axis and minor axis is?

OpenStudy (shamil98):

a^2 is ALWAYS the bigger denominator. 2a = major axis 2b = minor axis a^2 = 16 a =4 2a = major axis 2(4) = 8 major axis is 8 minor axis = 2b b^2 = 4 b = 2 minor axis = 4

OpenStudy (anonymous):

ahhh, I finally see what you did, thats a lot easier than I thought! how do I find the eccentricity? I know you had said it was like c/a, but what number represents c?

OpenStudy (shamil98):

Like in a triangle but the formula is. a^2 = b^2 + c^2 so 16 = 4 + c^2 12 = c^2 2√3 = c eccentricity = c/a 2√3/ 4 = √3/2

OpenStudy (anonymous):

thank you SOOO much! you are a life saver! How did you know that c^2 was 12 though?

OpenStudy (shamil98):

a^2 = b^2 + c^2 i just plugged in your values.

OpenStudy (shamil98):

a^2 = 16 b^2 = 4 and i solved for c :)

OpenStudy (anonymous):

oh okay! :)